Fynn performs an experiment where he flips a coin three times. If he performs this
experiment 100 times, how many repetitions of the experiment can he expect to result in the coin landing on the same side for all three flips?

Answer by greenestamps(13198)



P(all 3 heads) = (1/2)^3 = 1/8
P(all 3 tails) = (1/2)^3 = 1/8

P(all 3 the same) = 1/8 + 1/8 = 1/4

The expected number of times in 100 trials that the three coins land on the same side is 100(1/4) = 25.

ANSWER: 25